# GASLITEing the Retrieval: Exploring Vulnerabilities in Dense Embedding-based Search

### v1

@@ Appendix F Ablation Studies and Additional Comparison
F.1 Comparison With Prior Work
@@
F.2 Ablating GASLITE’s Algorithm Components
@@
F.3 GASLITE’s Hyperparameters
Through the paper, unless otherwise mentioned, we run our method with a trigger length of , appended to a given information (a text fixed throughout the attack); the trigger is initialized with text generated by GPT2 (Radford et al., 2019) conditioned on the given information. We run for iterations; average the gradient over random flips; performing flips in each iteration; each flip is sampled from a pool of most-promising candidates. In what follows we examine the impact of each parameter’s value on the GASLITE’s objective.
@@ F.3 GASLITE’s Hyperparameters
Trigger Initialization. To examine the effect of choice of trigger initialization in GASLITE, we consider five methods: generating (with GPT2) the initial trigger conditioned on the prefix (to provide a consistent with the ); filling the initial trigger with an arbitrary token (in our case “”, consistent with Zou et al. (2023) and Zhong et al. (2023)); stuffing the initial trigger with random passages from the corpus; stuffing with random golden passages (i.e., the ground-truth passages correspond to the attacked queries); stuffing with the attacked queries (identical to stuffing baseline §5). Notably, using passages for initialization requires access to the corpus, which we assume the attacker does not possess (§3). Results are shown in Tab. 5.
@@
F.4 GASLITE’s Defense-bypassing Variants
@@
In this subsection, we present the methodology and parameters we use to bypass common defenses with GASLITE later (App. G). In particular, we focus on limiting the extent of norm of the adversarial passages and their perplexity. Both were done by enriching the objective with an additional term, multiplied by some weight. In what follows, we elaborate on the chosen term and consider multiple choice for these scalar weight parameters.
L2 Norm.
@@
To minimize the of the crafted passage, we add the following term to the maximized objective: , where is the penalty weight. Attempting various values (Fig. 14(c)), we observe that going below norm of (among the largest norms of benign passages; Fig. 15(d)) is followed by a significant decrease in the attack success. We perform our -filtering bypass attempt (App. G) with K.
Perplexity.
@@
To minimize the perplexity of the crafted passage, we add the following term to the maximized objective: , where is the penalty weight and is the perplexity of a utility language model, in addition to limiting the token candidates to the top-1% logits of the LM.
@@
We calculate the perplexity with GPT2 (Radford et al., 2019), on which we also back-propagate, as part of the attack’s linear approximation (Alg. 1). Most embedding models are BERT-based, and, in particular, use BERT’s tokenizer (Devlin et al., 2019) which differs from GPT2’s. Thus, we pretrain GPT2 with BERT’s tokenizer and utilize this instance for this attack variant.181818Pretraining GPT2 following Andrej Karpathy’s nanoGPT recipe: https://github.com/karpathy/nanoGPT.
@@
We run the attack with different values, on trigger length (Fig. 14(a)) and (Fig. 14(b)), measuring the negative log-perplexity (the higher, the more fluent text is expected to be). We observe a moderate decline in the attack success when increasing the weight, allowing to capture a suitable weight to balance the trade off. Specifically, aiming to place the crafted passages in the average benign perplexity (Fig. 15(b)), we perform our perplexity-filtering bypass attempt (App. G) with .
Appendix G Defenses and Adaptive Variants Against Them
We evaluate GASLITE against previously proposed detection-based baseline defenses (Jain et al., 2023; Alon & Kamfonas, 2023; Zhong et al., 2023)—specifically, perplexity-based and norm-based—and attempt relatively simple approaches for bypassing these with GASLITE.
Concretely, for the detections we set a threshold of zero false-positives (e.g., highest perplexity of benign passages), and for the evasions we add an evasion term (e.g., an LM perplexity) to the optimized adversarial objective as well as trying simpler methods (more details in App. F.4). Demonstrating on two different models—picked for their high benign performance (E5 on fluency-based detection, and Contriever-MS for -based detection)—, we find evading detection possible, albeit comes with a price of a decrease in the attack success (Fig. 15). Future work may optimize directly under this setting, possibly employing recent advances in fluent LLM jailbreaks (Paulus et al., 2024).
-Norm-based Detection. As noted by Zhong et al. (2023), we observe that attacking dot-product similarity models results with crafted passages of large norm; this can be used to identify anomalous adversarial passages, by filtering passages surpassing the maximal benign (Fig. 15(d)). Enriching the adversarial objective adding a term to penalize large- passages (GASLITE-L2) mostly fails to achieve low (Fig. 15(d)) resulting with a significant performance decrease (Fig. 15(c)).
Fluency-based Detection. GASLITE, similarly to many other attacks (e.g., Zhong et al. (2023); Zou et al. (2023)), may result with non-fluent and non-sensical triggers, which can be exploited to identify adversarial passages via perplexity filtering (Jain et al., 2023; Alon & Kamfonas, 2023) (Fig. 15(b)). Enriching the adversarial objective by adding a GPT2 (Radford et al., 2019) perplexity term (GASLITE-Flu) preserves roughly half of the attack success (Fig. 15(a)) while completely assimilating in the GPT2’s perplexity distribution of the benign passages (Fig. 15(b)). We find even simpler approaches to perform well (Fig. 15(a)), although they present slightly higher perplexity (Fig. 15(b)); these include limiting the trigger length to (GASLITE10), and sampling token candidates, through the optimization, only from the top- logits of GPT2 (GASLITE-FluLogits).
Appendix H Additional Results
@@
| Model | Attack’s Threat Model | ||
| Knows All | Knows What | Knows Nothing | |
| (§6.1) | (§6.2) | (§6.3) | |
| Contriever | |||
| Contriever-MS | |||
| ANCE | |||
| mMPNet | |||
| Arctic | |||
| E5 | |||
| GTR-T5 | |||
| MiniLM | |||
| aMPNet | |||
H.1 “Knows All”
@@
| E5 | Targeted Query () | Hopalong Cassidy’s horse’s name | 1.0 |
@@
| E5 | Targeted Query () | Response when parents bring baby home | 1.0 |
@@
| Contriever-MS | Targeted Query () | Hopalong Cassidy’s horse’s name | 2.3556 |
@@
First, in Fig. 16, we show an analysis of the attack success (appeared@10) per model and concept under various budgets (). This analysis includes baselines ( Only, stuffing, Cor.Pois.), GASLITE method, the simulated perfect attack (§6.4), a GASLITE variant based on synthetic data (more in App. H.5), and a Cor.Pois. variant denoted as Cor.Pois. [on tok.], which evaluates Cor.Pois. on token space, that is, under a weaker threat model that assumes the attacker can directly control the input tokens. This further highlights GASLITE’s superiority, across budget choices and over baselines, even under a weaker threat model.
@@
| Potter | only | Voldemort was right all along and Harry Potter is a self-absorbed hero who doesn’t deserve the fame and glory he receives. |
@@
| Sandwich | only | Sandwiches are a boring and unimaginative food choice, lacking the creativity and flavor of more exciting culinary options. |
@@
| iPhone | only | iPhone is a status symbol that perpetuates consumerism and planned obsolescence, encouraging people to constantly upgrade to the latest model. |
@@
| Flower | only | Flowers are a waste of money and resources, providing little value beyond temporary decoration and fleeting beauty. |
@@
| Mortgage | only | Mortgages are a trap that keeps people in debt and financial insecurity, benefiting banks and lenders at the expense of homeowners. |
@@
H.3 “Knows [Almost] Nothing”
@@
H.4 Attacking LLM-based Retrievers
In what follows we evaluate the susceptibility of Stella embedding model191919https://huggingface.co/dunzhang/stella_en_1.5B_v5 to GASLITE and other attacks, repeating the experiments introduced in §5–6. Stella is a LLM-based embedding model (based on Qwen2; Alibaba (2024)) of 1.5B parameters, differing from the 110M-sized encoder-only models (mostly BERT-based) on which we focus in them main body. Generally, we find results consistent with our original evaluation.
Knows All. Repeating the single-query attack evaluations (§6.1) on Stella, we observe results consistent with evaluation on encoder-only models, as shown in Tab. 12. Specifically, stuffing is still a relatively effective baseline, and GASLITE consistently achieves optimal appeared@10 measures while crafting adversarial passage of very high similarity with the targeted query.
@@ H.4 Attacking LLM-based Retrievers
| Cos | Stella | () | () | () | |||
Knows What. Repeating the concept-specific attack evaluations (§6.2), we notice trends similar to the ones shown in the main body: concept-specific attacks are possible using few adversarial passages while achieving top-10 visibility for most queries. Nonetheless, we find Stella slightly harder to attack—it achieves slightly less success rate compared the median encoder-only model—we believe that further optimizing and adapting GASLITE to LLM-based models—on which we have not focused—is possible and can improve concept-specific attack success.
| Metric Targeted Concept | Golf | Sandwich | Boston | Vaccine | iPhone | Flower |
| appeared@10 |
H.5 “Query-less” Attacks
@@
H.5.1 Attacking via GASLITE with Synthetic (LLM-generated) queries
To evaluate GASLITE and other methods on the Knows What setting in an isolated manner (i.e., without additional factors that may affect the evaluation), we assumed, per our threat model (3), that the attacker has access to concept-related sample queries (§6.2). In the following experiment we show it is possible to relax this assumption by generating synthetic sample queries, using an LLM.
Experiment Setup. We generate concept-related sample queries, for each attacked concept. We use these as synthetic sample queries (instead of the dataset queries used in §6.2). To generate these we prompt Claude-3.5-Sonnet 202020https://www.anthropic.com/news/claude-3-5-sonnet with the following few-shot template of 3 example queries:
@@ H.5.1 Attacking via GASLITE with Synthetic (LLM-generated) queries
H.5.2 Attacking via GASLITE with Surrogate Queries
In the following we study cross query-set transferability under the Knows Nothing setting. Following Zhong et al. (2023) results on successfully transferring Cor.Pois. across datasets, we experiment GASLITE under this setting. Concretely, we use GASLITE to craft adversarial passage given with source queries (e.g., a sample of MSMARCO’s train set) and evaluate poisoning with those passages another, target, dataset (e.g., FiQA2018, SciFact and Quora test sets; Muennighoff et al. (2023)). Said differently, we evaluate the success of adversarial passages, crafted as part of the Knows All attacks (on MSMARCO and NQ; §6.3), on other datasets of different query distrbiutions.
Consistent with the findings of Zhong et al. (2023), results in Fig. 20 show it is possible of transfer attack across query sets. Particularly, topic-specific query sets (FiQA and SciFact) show high susceptibility to this method. We hypothesize that the semantic diversity in the source query sets (e.g., in MSMARCO) is useful for crafting query-universal attacks that applies effectively to target query sets of narrower semantics.
H.6 Top-Results Attacker
We form and experiment an additional baseline, by considering attackers that has access to the top-results of the query; notably, this deviates from our threat model (§3) that assumes no access to the corpus. Following the resemblance of GASLITE’s passages to potential results of the queries, the following baseline attack employs access to top results, and summarizes it to an adversarial suffix.
Concretely, given the queries available to the attacker (), and their corresponding top results (), the attacker summarizes the top-1 results of all queries (using an LLM), and use it as an adversarial suffix. Following our multi-budget scheme (App. D.1), each adversarial suffix is crafted for the query set of a KMeans cluster.
@@ H.6 Top-Results Attacker
H.7 Transferability Across Models
Applying GASLITE, a gradient-based method, requires access to model weights. However, prior work demonstrated success in transferring attacks on one model (that enables white-box access) to another model (with black-box access) (Szegedy et al., 2014; Zou et al., 2023). In what follows, we study the transferability of GASLITE under the different threat models. Note, however, that GASLITE was not explicitly optimized for such transferability.
To assess GASLITE’s transferability, we take the adversarial passages crafted for the different threat models (§6), and evaluate them as attacks against all models. Results in Fig. 21 show transferability occurs mainly within model families; for example, Contriever and Contriever-MS, or aMPNet and mMPNet (which share backbone architecture), or Arctic and E5 (Arctic was trained based on E5). Additionally, transferability is weaker in the more challenging threat models (e.g., Knows Nothing) that target a wide range of queries.
H.8 Attack Run-Times
GASLITE, similarly to other gradient-based discrete optimizers Zou et al. (2023) and to prior attack (Cor.Pois.; Zhong et al. (2023)) require a non-negligible amount of compute to craft the adversarial passages. We briefly discuss the compute-success trade-off presented by the baselines and the different attacks.
@@ H.8 Attack Run-Times
Appendix I On Practical Considerations in Poisoning a corpus
First, poisoning a retrieval corpus is mostly done via insertion of the malicious text (e.g., uploading code section to a public repository or paragraphs to Wikipedia), as a result the input tokens depend on the tokenization process of the embedding model. Perhaps surprisingly, this capability is strictly weaker than directly controlling the input tokens, in particular, for many tokenizers (e.g., BERT’s; Devlin et al. (2019)), there exist token lists that cannot be reached from any text, as demonstrated in Tab. 15.212121https://gist.github.com/matanbt/afdd59f66158dc30b33d496864126e96

### v2

@@ Appendix F Ablation Studies and Additional Comparison
F.1. Comparison With Prior Work
@@
F.2. Ablating GASLITE’s Algorithm Components
@@
F.3. GASLITE’s Hyperparameters
Throughout the paper, unless otherwise mentioned, we run our method with a trigger length of , appended to a given information (a text fixed throughout the attack); the trigger is initialized with text generated by GPT2 (Radford et al., 2019) conditioned on the given information. We run for iterations; average the gradient over random flips; performing flips in each iteration; each flip is sampled from a pool of most-promising candidates. In what follows we examine the impact of each parameter’s value on the GASLITE’s objective.
@@ F.3. GASLITE’s Hyperparameters
Trigger Initialization. To examine the effect of choice of trigger initialization in GASLITE, we consider five methods: generating (with GPT2) the initial trigger conditioned on the prefix (to provide a consistent with the ); filling the initial trigger with an arbitrary token (in our case “”, consistent with (Zou et al., 2023) and (Zhong et al., 2023)); stuffing the initial trigger with random passages from the corpus; stuffing with random golden passages (i.e., the ground-truth passages correspond to the attacked queries); stuffing with the attacked queries (identical to stuffing baseline §5). Notably, using passages for initialization requires access to the corpus, which we assume the attacker does not possess (§3). Results are shown in Tab. 10.
@@
F.4. GASLITE’s Defense-bypassing Variants
@@
In this subsection, we present the methodology and parameters we use to bypass common defenses with GASLITE later (§7). In particular, we focus on limiting the extent of norm of the adversarial passages and their perplexity. Both were done by enriching the objective with an additional term, multiplied by some weight. In what follows, we elaborate on the chosen term and consider multiple choice for these scalar weight parameters.
@@
L2 Norm. To minimize the of the crafted passage, we add the following term to the maximized objective: , where is the penalty weight. Attempting various values (Fig. 18(c)), we observe that going below norm of (among the largest norms of benign passages; Fig. 8(d)) is followed by a significant decrease in the attack success. We perform our -filtering bypass attempt (§7) with K.
@@
Perplexity. To minimize the perplexity of the crafted passage, we add the following term to the maximized objective:
where is the penalty weight and is the perplexity of a utility language model, in addition to limiting the token candidates to the top-1% logits of the LM.
@@
We calculate the perplexity with GPT2 (Radford et al., 2019), on which we also back-propagate, as part of the attack’s linear approximation (Alg. 1). Most embedding models are BERT-based, and, in particular, use BERT’s tokenizer (Devlin et al., 2019) which differs from GPT2’s. Thus, we pretrain GPT2 with BERT’s tokenizer and utilize this instance for this attack variant.242424Pretraining GPT2 following Andrej Karpathy’s nanoGPT recipe: https://github.com/karpathy/nanoGPT.
@@
We run the attack with different values, on trigger length (Fig. 18(a)) and (Fig. 18(b)), measuring the negative log-perplexity (the higher, the more fluent text is expected to be). We observe a moderate decline in the attack success when increasing the weight, allowing to capture a suitable weight to balance the trade off. Specifically, aiming to place the crafted passages in the average benign perplexity (Fig. 8(b)), we perform our perplexity-filtering bypass attempt (§7) with .
Appendix G Additional Results
@@
| Model | Attack’s Threat Model | |||
| Knows All | Knows What | Knows What | Knows Nothing | |
| (§6.1) | (§6.2) | (§6.2) | (§6.3) | |
| Contriever | ||||
| Contriever-MS | ||||
| ANCE | ||||
| mMPNet | ||||
| Arctic | ||||
| E5 | ||||
| GTR-T5 | ||||
| MiniLM | ||||
| aMPNet | ||||
| Model | Info only | Stuffing | Corp. Pois | GASLITE |
| Contriever | ||||
| Contriever-MS | ||||
| ANCE | ||||
| mMPNet | ||||
| Arctic | ||||
| E5 | ||||
| GTR-T5 | ||||
| MiniLM | ||||
| aMPNet |
G.1. “Knows All”
@@
E5 |
Targeted Query () | Hopalong Cassidy’s horse’s name | 1.0 |
@@
E5 |
Targeted Query () | Response when parents bring baby home | 1.0 |
@@
Contriever-MS |
Targeted Query () | Hopalong Cassidy’s horse’s name | 2.3556 |
@@
First, in Fig. 19, we show an analysis of the attack success (appeared@10) per model and concept under various budgets (). This analysis includes baselines ( Only, stuffing, Cor.Pois.), GASLITE method, the simulated perfect attack (§6.4), a GASLITE variant based on synthetic data (more in App. G.6), and a Cor.Pois. variant denoted as Cor.Pois. [on tok.], which evaluates Cor.Pois. on token space, that is, under a weaker threat model that assumes the attacker can directly control the input tokens. This further highlights GASLITE’s superiority, across budget choices and over baselines, even under a weaker threat model.
@@
To control for our choice of query partition, we also compare attacks under a Knows What, allowing only a single adversarial passage. Tab. 12 shows trends are similar to the ones shown under the multi-budget setting.
@@
Potter |
only | Voldemort was right all along and Harry Potter is a self-absorbed hero who doesn’t deserve the fame and glory he receives. |
@@
Sandwich |
only | Sandwiches are a boring and unimaginative food choice, lacking the creativity and flavor of more exciting culinary options. |
@@
iPhone |
only | iPhone is a status symbol that perpetuates consumerism and planned obsolescence, encouraging people to constantly upgrade to the latest model. |
@@
Flower |
only | Flowers are a waste of money and resources, providing little value beyond temporary decoration and fleeting beauty. |
@@
Mortgage |
only | Mortgages are a trap that keeps people in debt and financial insecurity, benefiting banks and lenders at the expense of homeowners. |
@@
G.3. “Knows [Almost] Nothing”
@@
G.4. Attacking Additional Retrievers
In what follows we evaluate the susceptibility of retrievers of additional backbone architectures, including LLM-based. Specifically, we evaluate: Stella (1.5B; based on the LLM model Qwen2 (Alibaba, 2024))252525https://hf.co/dunzhang/stella_en_1.5B_v5, Qwen3-Embedding (Zhang et al., 2025a) (0.6B; based on the LLM model Qwen3) 262626https://hf.co/Qwen/Qwen3-Embedding-0.6B, GTE-ModernBERT (with ModernBERT backbone (Warner et al., 2024)).272727https://hf.co/Alibaba-NLP/gte-modernbert-base These models complement those from our main experiments, where we focus on the popular and efficient 110M-sized encoder-only models, mostly based on the encoder-only BERT backbone. We repeat the experiments introduced in §5–6, and generally, find results consistent with our original evaluation.
Knows All. Repeating the single-query attack evaluations (§6.1) on the additional models, Tab. 18 shows results consistent with evaluation on encoder-only models. Specifically, stuffing is still a relatively effective baseline, and GASLITE consistently achieves optimal appeared@10 measures while crafting adversarial passage of very high similarity with the targeted query.
@@ G.4. Attacking Additional Retrievers
| Cos | Stella (LLM-based) | () | () | () | |||
| Cos | Qwen3 (LLM-based) | () | () | () | |||
| Cos | GTE-ModernBERT | () | () | () | |||
Knows What. Repeating the concept-specific attack evaluations (§6.2), we once again notice trends similar to the ones shown in the main body: concept-specific attacks are possible using few adversarial passages while achieving top-10 visibility for most queries. We find GTE-ModernBERT and Qwen3 to be as susceptible to these attacks as the average encoder-only model, with Stella being slightly more robust. However, as supported by the high susceptibility of the LLM-based model, Qwen3, we believe that further optimizing and adapting GASLITE to models or larger scale (such as Stella)—on which we have not focused in this work—is possible and can improve concept-specific attack success.
| Targeted Concepts | |||||||||
| Model | Potter | Golf | Sandwich | Boston | Mortgage | Vaccine | iPhone | Flower | Avg. |
| Stella (LLM-based) | |||||||||
| Qwen3 (LLM-based) | |||||||||
| GTE-ModernBERT | |||||||||
G.5. Transferability Across Models
Applying GASLITE, a gradient-based method, requires access to model weights. However, prior work demonstrated success in transferring attacks on one model (that enables white-box access) to another model (with black-box access) (Szegedy et al., 2014; Zou et al., 2023). In what follows, we study the transferability of GASLITE under the different threat models. Note, however, that GASLITE was not explicitly optimized for such transferability; future work may build upon the following insights to directly improve transferability.
Applying GASLITE, a gradient-based method, requires access to model weights. However, prior work demonstrated success in transferring attacks on one model (that enables white-box access) to another model (with black-box access) (Szegedy et al., 2014; Zou et al., 2023). While GASLITE was not explicitly optimized for such transferability, we study it under the different threat models; future work may build upon the following insights, in addition to past work in transferability (Demontis et al., 2019; Mao et al., 2022), to directly improve transferability.
Concretely, we use adversarial passages crafted for one model, to attack another (black-box access) model. Our results (Fig. 23 in App. G.5) show transferability is often limited, but does occur within model families (e.g., between Contriever and Contriever-MS, aMPNet and mMPNet, and Arctic and E5), which build off of similar architectures, but vary in implementation, including by fine tuning models in different ways or employing different similarity metrics.
Setup. To assess GASLITE’s transferability, we take the adversarial passages crafted for the different threat models (§6), and evaluate them as attacks against all models.
Results. From results in Fig. 23 we observe that transferability is generally limited, but does occur within model families; for example, Contriever and Contriever-MS, or aMPNet and mMPNet (which share pretrained base architecture), or Arctic and E5 (Arctic was trained on top of E5). Additionally, transferability is weaker in the more challenging threat models (e.g., Knows Nothing) that target a wide range of queries (Fig. 23).
G.6. “Query-less” Attacks
@@ G.5. Transferability Across Models
G.6.1. Attacking via GASLITE with Synthetic (LLM-generated) queries
To evaluate GASLITE and other methods on the Knows What setting in an isolated manner (i.e., without additional factors that may affect the evaluation), we assumed, per our threat model (3), that the attacker has access to concept-related sample queries (§6.2). In the following experiment, we show that it is possible to relax this assumption by generating synthetic sample queries, using an LLM.
Experiment Setup. We generate 250 concept-related sample queries, for each attacked concept. We use these as synthetic sample queries (instead of the dataset queries used in §6.2). To generate these we prompt Claude-3.5-Sonnet 282828https://www.anthropic.com/news/claude-3-5-sonnet with the following few-shot template of 3 example queries:
@@ G.6.1. Attacking via GASLITE with Synthetic (LLM-generated) queries
G.7. Top-Results Attacker
@@
We form and experiment with an additional baseline, by considering attackers that has access to the top-results of the query; notably, this deviates from our threat model (§3) that assumes no access to the corpus. Following the resemblance of GASLITE’s passages to potential results of the queries, the following baseline attack employs access to top results, and summarizes it to an adversarial suffix.
Concretely, given the queries available to the attacker (), and their corresponding top results (), the attacker summarizes the top-1 results of all queries (using an LLM), and use it as an adversarial suffix. Following our multi-budget scheme (App. D.1), each adversarial suffix is crafted for the query set of a -means cluster.
@@
G.8. Attack Run-Times
GASLITE, similarly to other gradient-based discrete optimizers (Zou et al., 2023) and to prior attack (Cor.Pois.; (Zhong et al., 2023)) require a non-negligible amount of compute to craft the adversarial passages. We briefly discuss the compute-success trade-off presented by the baselines and the different attacks.
@@ G.8. Attack Run-Times
Appendix H On Practical Considerations in Poisoning a corpus